OpenStudy (anonymous):

-5/6n=40 solve (Points : 1) 56 48 –48 –56

myininaya (myininaya):

\[\frac{-5}{6}n=40 ?\]

OpenStudy (anonymous):

yes thats it

myininaya (myininaya):

Try to multiply both sides by the reciprocal on -5/6 on both sides

myininaya (myininaya):

\[\frac{-5}{6}n=40 \\ \frac{\cancel{-6}}{\cancel{5}} \cdot \frac{\cancel{5}}{\cancel{-6}}n=\frac{-6}{5} \cdot 40\]

OpenStudy (anonymous):

i know that but i got confused when it says multiply - \[\frac{ 5}{ 6 } \] by 40

OpenStudy (anonymous):

and i used my calculator but that was wrong

myininaya (myininaya):

so you multiplied 40 and -6/5 ?

myininaya (myininaya):

Because that is what it says above.

OpenStudy (anonymous):

ok i got -33.3333333333333